OpenStudy (anonymous):

which equation represents the line passing through the points(3,2)and(-9,6)

OpenStudy (anonymous):

we will write it in y=mx +b form okay? So first we need to find the slope(m) the equation is y2-y1 OVER(/) x2-x1 . So plug in the values; (6)-(2)/(-9)-(3) which is equal to -1/3 . so know we have to use point slope form. Y-y1=m(x-x1) and plug in only one point(you can choose any but lets chose the first one) Y-2=-1/3(x-3) which is y=-1/3x +3